A circle is tangent to the y axis at (0,7) and the center is on the line y = 4x-9. Give the equation of the line, in polynomial form.
Answer by lwsshak3(11628)

A circle is tangent to the y axis at (0,7) and the center is on the line y = 4x-9. Give the equation of the line, in polynomial form.
***
The line y=7 intersects with line y=4x-9
7=4x-9
4x=16
x=4
center: (4,7)
r=4
Equation: (x-4)^2+(y-7)^2=16
in polynomial form:
x^2-8x+16+y^2-14y+49=16
x^2+y^2-8x-14y=-49